# Towards Adaptive Visual Token Pruning for Large Multimodal Models

### v1

@@ Abstract
Introduction
@@
Large Multimodal Models (LMMs) (Bai et al. 2025; Team et al. 2025; Zhu et al. 2025; Liu et al. 2024a; Lin et al. 2023) have substantially enhanced the reasoning capabilities of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023; Chiang et al. 2023; Zhu et al. 2023; Li et al. 2023a; Zhang et al. 2023; Huang et al. 2023; Wang et al. 2023) by enabling joint processing of multimodal inputs such as images and texts. Typically, visual inputs are encoded into dense token sequences via a vision encoder and concatenated with textual tokens for unified processing by the language model. However, the resulting token sequences often reach thousands in length, leading to significant computational and memory overhead due to the quadratic complexity of self-attention with respect to sequence length (Vaswani et al. 2017; Face 2024; Chen et al. 2023; Keles, Wijewardena, and Hegde 2023; Liu et al. 2022). These limitations present major obstacles to deploying LMMs in resource-constrained or latency-sensitive environments (Chen et al. 2024a; Lin et al. 2025a).
@@
Recent studies have shown that visual token representations in LMMs exhibit substantial redundancy (Liu et al. 2024c; Shang et al. 2024a; Huang et al. 2024; Tong et al. 2025; Li, Yang, and Lu 2025). Leveraging this insight, visual token pruning methods have been proposed to reduce computational cost by selectively removing unnecessary tokens. By eliminating redundant visual tokens, these approaches effectively alleviate the quadratic burden of long input sequences. Notably, prior work demonstrates that pruning the majority of visual tokens can yield significant efficiency gains with minimal performance degradation (Zhang et al. 2024a; Chen et al. 2024a; Lin et al. 2025a; Huang et al. 2024; Sun et al. ).
@@
While token pruning is beneficial, its application to LMMs remains challenging. Existing token pruning methods can be broadly categorized into three types. 1) A common approach is to use attention scores to identify redundant tokens (Chen et al. 2024b; Lin et al. 2025b; Shang et al. 2024b; Tong et al. 2025). However, such methods are susceptible to positional bias and often retain spatially adjacent tokens with high similarity, leading to redundancy and performance degradation. 2) Other methods rely on model-specific calibration or fine-tuning (Ye et al. 2025; Lin et al. 2025b; Li et al. 2025b; Cai et al. 2024), which incurs high computational cost and limits scalability in practical deployment. 3) Another line of work addresses token pruning by maximizing the minimum distance between tokens (Alvar et al. 2025). While this strategy encourages token separation, it is sensitive to outliers and may fail to preserve cross-modal alignment, ultimately affecting downstream performance.
@@
Related Work
@@
Large Multimodal Models (LMMs) extend pretrained LLMs by integrating multiple modalities (Bai et al. 2025; Team et al. 2025), such as images and texts. Typically, a vision encoder extracts dense visual features, which are projected into the language model’s embedding space via modules like Q-Former (Chebotar et al. 2023) or MLPs (Taud and Mas 2017). Conventional methods resize high-resolution images to a fixed scale (Koonce 2021), causing geometric distortions and loss of fine-grained spatial details. Dynamic tiling mitigates this by splitting images into smaller regions, each independently encoded by a shared encoder, better preserving local information (Yuan et al. 2021; Yin et al. 2022). However, this increases the number of visual tokens and computational cost, a challenge further amplified in video LMMs due to multi-frame processing (Liu et al. 2024c). These challenges highlight the need for efficient inference methods to enable LMMs deployment under resource constraints (Tong et al. 2025).
@@
Visual Token Pruning aims to reduce computational overhead and improve inference efficiency by removing redundant or less informative visual tokens. A common strategy is to leverage attention scores to identify tokens for removal. For example, PruMerge (Shang et al. 2024b) clusters and merges tokens in the vision encoder based on attention sparsity, while FastV (Chen et al. 2024a) utilizes attention weights from the second layer of the LLM to guide pruning. SparseVLM (Zhang et al. 2024b) employs cross-modal attention for text-conditioned token selection, and VisionZip (Yang et al. 2025) compresses visual inputs via CLS token attention in the final vision encoder layer. FlowCut (Tong et al. 2025) identifies redundancy from an information flow perspective by analyzing attention propagation. LVPruning (Sun et al. 2025) uses cross-attention to assess vision token importance via interaction with language tokens, guiding token pruning. Other methods adopt calibration-based strategies, where pruning ratios and depths are determined by evaluating model behavior on a held-out set. FitPrune (Ye et al. 2025) compares attention distributions before and after pruning to inform token selection, while VTW (Lin et al. 2025a) shows that tokens can be safely dropped after specific layers, guided by calibration. In addition, approaches like DivPrune (Alvar et al. 2025) addresses token pruning by maximizing the minimum distance between tokens. Moreover, a concurrent work (Zhang et al. 2025) reformulates the token pruning problem using determinantal point processes to achieve effective pruning.
Figure 1: Overview of the visual token pruning method. Visual tokens that exhibit low semantic alignment with textual tokens are pruned by measuring their cross-modal mutual information. The remaining tokens are further reduced by maximizing the intra-modal representation within the visual modality, which is efficiently solved using a greedy algorithm.
Token Redundancy Across Modalities
@@
Previous studies have yet to systematically and thoroughly compare token redundancy across different modalities (Shang et al. 2024a; Alvar et al. 2025; Li, Yang, and Lu 2025). In this section, we analyze and contrast the redundancy of textual and visual tokens from three perspectives: attention mechanisms, semantic distribution and information repetition. Our findings reveal that visual tokens exhibit substantially higher redundancy than textual tokens. Motivated by this observation, we retain all textual tokens during the token pruning process and apply pruning exclusively to visual tokens.
The attention distribution in multimodal models is highly skewed, with textual tokens receiving significantly higher weights than visual tokens (Li et al. 2025a; Lee et al. 2025). This suggests that the model primarily relies on text for semantic understanding. Visual inputs often consist of hundreds or thousands of tokens, far exceeding the number of textual tokens. Given the quadratic complexity of attention with respect to token count, this mismatch between quantity and contribution is pronounced. Despite the large number of visual tokens, their impact on the final representation is limited, resulting in poor trade offs between computational cost and semantic gain and reflecting the high redundancy of visual tokens.
Each token in text typically carries explicit semantic meaning, such as a noun, verb, or conjunction, reflecting natural language as a highly optimized discrete encoding system with dense and relatively uniform information distribution (Hirschberg and Manning 2015). In contrast, semantic information in images is highly concentrated in a small number of salient regions, such as foreground objects, while the majority of visual tokens correspond to low-semantic areas like sky or walls, which mostly contain low-frequency textures or superficial variations and contribute little to high-level semantic understanding. Consequently, the effective semantic density of visual tokens is significantly lower than that of textual tokens.
Compared to textual tokens, visual tokens exhibit higher information redundancy due to differences in their generation processes and underlying information structures. Textual tokens, derived from natural language, follow Zipf’s law (Newman 2005; Adamic and Huberman 2002) and represent distinct semantic units with low repetition. In contrast, visual tokens are typically generated by uniformly partitioning images or extracting low-level features without semantic organization. For example, a sky region is divided into multiple similar patches that cluster in embedding space and jointly represent a single high-level concept, causing redundancy. Visual tokenization resembles physical partitioning rather than frequency-driven semantic abstraction, forcing models to process many similar tokens and resulting in computational and representational inefficiency.
Methodology
Problem Formulation
@@
Our pruning pipeline can be roughly divided into two stages. In the first stage, we select a subset $\mathcal{X}^{v(1)}\subset\mathcal{X}^{v}$ with size $\left|\mathcal{X}^{v(1)}\right|={N}_{1}$, where ${N}_{1}<N$, by retaining tokens with the highest semantic alignment to the text modality, thereby ensuring the preservation of tokens with strong cross-modal correlations. In the second stage, we further prune $\mathcal{X}^{v(1)}$ to obtain a more compact subset $\mathcal{X}^{v(2)}\subset\mathcal{X}^{v(1)}$ with size $\left|\mathcal{X}^{v(2)}\right|={N}_{2}$, where ${N}_{2}<{N}_{1}$, by maximizing the expected pairwise distance in the embedding space. This promotes non-redundant and information-rich representations while reducing computational overhead. Figure 1 illustrates the overall architecture of our method.
Cross-Modal Alignment Aware Token Filtering
During training, multimodal models primarily optimize cross-entropy loss for text generation, which often leads to an implicit weakening of cross-modal alignment (Covert et al. 2024). Pruning as a post-processing step for the model is crucial for cross-modal alignment. To preserve cross-modal alignment during visual token pruning, we introduce a mutual information based criterion that selects a subset of visual tokens $\mathcal{X}^{v(1)}$ which maximizes the expected mutual information with the textual tokens. Our selection objective can be formulated as follows:
$\mathcal{X}^{v(1)}=\mathop{\arg\max}_{\begin{subarray}{c}\mathcal{X}^{v^{\prime}}\subset\mathcal{X}^{v},|\mathcal{X}^{v^{\prime}}|=N_{1}\end{subarray}}\mathbb{E}_{x^{t}\in\mathcal{X}^{t}}\left[I(\mathcal{X}^{v^{\prime}};x^{t})\right]$ (1)
@@
This score can be interpreted as the inverse of the InfoNCE loss  (Rusak et al. 2024), reflecting the average semantic alignment strength of visual tokens within the shared embedding space. Intuitively, tokens with higher alignment scores are retained, while those with lower semantic alignment are pruned.
Although mutual information provides a theoretically grounded measure of cross-modal alignment, its precise estimation in high dimensional spaces often requires costly density modeling. We approximate mutual information using the $L_{2}$ norm, which reduces computational overhead without compromising performance. It is worth noting that our approximation is grounded in mathematical theory and does not rely on any specific model architecture. Additionally, we provide theoretical justification and experimental analysis to explain why the $L_{2}$ norm is used for approximation instead of other metrics (see Appendix 1 and 2). Therefore, we can compute $\alpha_{i}$ using the following formula:
@@
$\mathcal{X}^{v(1)}=\left\{{x}_{i}^{v}\in\mathcal{X}^{v}\mid\alpha_{i}\in top_{N_{1}}(\alpha_{1},\alpha_{2},\ldots,\alpha_{N})\right\}$ (4)
This selection effectively filters out semantically misaligned visual tokens while retaining those with the strongest cross-modal correlations, thereby enhancing alignment quality and reducing computational overhead.
Intra-Modal Representational Maximization
@@
$\mathcal{X}^{v(2)}=\mathop{\arg\max}_{\begin{subarray}{c}\widetilde{\mathcal{X}}\subset\mathcal{X}^{v(1)},|\widetilde{\mathcal{X}}|=N_{2}\end{subarray}}\mathbb{E}_{\begin{subarray}{c}{x}_{i}^{v},{x}_{j}^{v}\in\widetilde{\mathcal{X}},i\neq j\end{subarray}}\left[D({x}_{i}^{v},{x}_{j}^{v})\right]$ (6)
@@
$\max_{\boldsymbol{\gamma}\in\{0,1\}^{N_{1}}}\sum_{i=1}^{N_{1}}\sum_{\begin{subarray}{c}j=1,j\neq i\end{subarray}}^{N_{1}}\gamma_{i}\gamma_{j}\cdot D({x}_{i}^{v},{x}_{j}^{v})\\ ,\quad\sum_{i=1}^{N_{1}}\gamma_{i}=N_{2}$ (7)
@@
At the initial stage of the greedy algorithm, we quantify pairwise relationships among visual tokens by constructing a cosine similarity matrix $\mathcal{C}\in\mathbb{R}^{N_{1}\times N_{1}}$. Each element $\mathcal{C}_{ij}$ of the matrix $\mathcal{C}$ represents the similarity between tokens $x_{i}^{v}$ and $x_{j}^{v}$. Based on this matrix, we compute the average similarity of each token with all others. The token with the lowest average similarity is chosen as the initial seed to initialize the selected set for the first iteration $\mathcal{S}^{(1)}$, and the remaining tokens constitute the initial remaining set $\mathcal{R}^{(1)}$. This can be formulated as follows:
@@
$\mathcal{S}^{(1)}=\{s^{(1)}\}\quad\mathcal{R}^{(1)}=[N_{1}]\setminus\{s^{(1)}\}$ (11)
@@
$\mathcal{S}^{(t)}=\mathcal{S}^{(t-1)}\cup\{s^{(t)}\},\quad\mathcal{R}^{(t)}=\mathcal{R}^{(t-1)}\setminus\{s^{(t)}\}$ (15)
@@
Experiments
Experimental Setup
Models and Baselines. We evaluate the performance of our method across several representative LMMs, including LLaVA-1.5-7B (Liu et al. 2023), LLaVA-1.5-13B (Liu et al. 2023), LLaVA-NEXT-7B (Liu et al. 2024b), LLaVA-NEXT-Video-7B (Zhang et al. 2024c) and Qwen2-VL-7B (Wang et al. 2024). These models span a range of parameter scales and encompass both image and video understanding tasks. We evaluate our method alongside several widely adopted baselines, including ToMe (Bolya et al. 2023), FastV (Chen et al. 2024a), SparseVLM (SpVLM) (Zhang et al. 2024b), PDrop (Xing et al. 2024) and DivPrune (Alvar et al. 2025). We additionally include the VTW (Lin et al. 2025b) method in our efficiency evaluation. All baselines are assessed under consistent experimental settings.
Datasets and Metrics. We evaluate our method on diverse multimodal benchmarks covering various tasks. For ChartQA (Masry et al. 2022), we report the relaxed score. Image captioning quality and diversity are measured by CIDEr on COCO (Sharma et al. 2018) and NoCaps (Agrawal et al. 2019). MMBenchEN (MMB) (Liu et al. 2024d) uses a GPT-based score. Perceptual understanding is assessed via the perception score on MME (Fu et al. 2023). Accuracy evaluates general reasoning on MMU (Zheng et al. 2025) and text recognition on OCRBench (Liu et al. 2024e). Exact match accuracy measures question answering on VQAOK (Marino et al. 2019) and VQATEXT (Singh et al. 2019). POPE (Li et al. 2023b) is evaluated with precision for positional understanding. This comprehensive protocol ensures thorough assessment of generalization across multimodal tasks.
Implementation Details. Our experiments are conducted using the PyTorch framework (Paszke et al. 2019) and the Hugging Face Transformers library (Wolf et al. 2020). We utilize an NVIDIA H100 GPU with 80GB of memory. Moreover, we use scikit-learn (Pedregosa et al. 2011) to compute mutual information, where the library employs a nonparametric $k$ nearest neighbors approximation method, with $k$ set to 3 by default. We set the parameter $N_{1}=\left\lfloor 0.8N\right\rceil$ by default. In both the ablation study and case study, we fix the final number of retained tokens to 64.

### v2

# TrimTokenator: Towards Adaptive Visual Token Pruning for Large Multimodal Models

@@ Abstract
TrimTokenator: Towards Adaptive Visual Token Pruning for Large Multimodal Models
Hao Zhang1thanks: These authors contribute equally to this work., Mengsi Lyu [^1] footnotemark: 1, Chenrui He1, Yulong Ao1thanks: Corresponding author., Yonghua Lin122footnotemark: 2 1Beijing Academy of Artificial Intelligence (BAAI)
1 Introduction
@@
Large Multimodal Models (LMMs) Bai et al. (2025); Team et al. (2025); Zhu et al. (2025); Liu et al. (2024a); Lin et al. (2023) have substantially enhanced the reasoning capabilities of Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chiang et al. (2023); Zhu et al. (2023); Li et al. (2023a); Zhang et al. (2023); Huang et al. (2023); Wang et al. (2023) by enabling joint processing of multimodal inputs such as images and texts. Typically, visual inputs are encoded into dense token sequences via a vision encoder and concatenated with textual tokens for unified processing by the language model. However, the resulting token sequences often reach thousands in length, leading to significant computational and memory overhead due to the quadratic complexity of self-attention with respect to sequence length Vaswani et al. (2017); Face (2024); Chen et al. (2023); Keles et al. (2023); Liu et al. (2022). These limitations present major obstacles to deploying LMMs in resource-constrained or latency-sensitive environments Chen et al. (2024a); Lin et al. (2025a).
@@
Recent studies have shown that visual token representations in LMMs exhibit substantial redundancy Liu et al. (2024c); Shang et al. (2024a); Huang et al. (2024); Tong et al. (2025); Li et al. (2025a). Leveraging this insight, visual token pruning methods have been proposed to reduce computational cost by selectively removing unnecessary tokens. By eliminating redundant visual tokens, these approaches effectively alleviate the quadratic burden of long input sequences. Notably, prior work demonstrates that pruning the majority of visual tokens can yield significant efficiency gains with minimal performance degradation Zhang et al. (2024a); Chen et al. (2024a); Lin et al. (2025a); Huang et al. (2024); Sun et al. .
@@
While token pruning is beneficial, its application to LMMs remains challenging. Existing token pruning methods can be broadly categorized into three types. 1) A common approach is to use attention scores to identify redundant tokens Chen et al. (2024b); Lin et al. (2025b); Shang et al. (2024b); Tong et al. (2025). However, such methods are susceptible to positional bias and often retain spatially adjacent tokens with high similarity, leading to redundancy and performance degradation. 2) Other methods rely on model-specific calibration or fine-tuning Lin et al. (2025b); Li et al. (2025c); Cai et al. (2024), which incurs high computational cost and limits scalability in practical deployment. 3) Another line of work addresses token pruning by maximizing the minimum distance between tokens Alvar et al. (2025). While this strategy encourages token separation, it is sensitive to outliers and may fail to preserve cross-modal alignment, ultimately affecting downstream performance.
Figure 1: Overview of the visual token pruning method. We compute the average mutual information (MI) between each visual token and textual tokens to obtain its semantic alignment score $\alpha_{i}$, and preserve the highest score tokens to form $\mathcal{X}^{v(1)}$. This subset is further refined by Greedy RepMax, which prunes redundant tokens to yield $\mathcal{X}^{v(2)}$. Greedy RepMax is a greedy approximation to the NP Hard problem of maximizing the expected pairwise distance among visual tokens.
@@
2 Related Work
2.1 Large Multimodal Models (LMMs)
@@
Large Multimodal Models (LMMs) extend pretrained LLMs by integrating multiple modalities Bai et al. (2025); Team et al. (2025), such as images and texts. Typically, a vision encoder extracts dense visual features, which are projected into the language model’s embedding space via modules like Q-Former Chebotar et al. (2023) or MLPs Taud and Mas (2017). Conventional methods resize high-resolution images to a fixed scale Koonce (2021), causing geometric distortions and loss of fine-grained spatial details. Dynamic tiling mitigates this by splitting images into smaller regions, each independently encoded by a shared encoder, better preserving local information Yuan et al. (2021); Yin et al. (2022). However, this increases the number of visual tokens and computational cost, a challenge further amplified in video LMMs due to multi-frame processing Liu et al. (2024c). These challenges highlight the need for efficient inference methods to enable LMMs deployment under resource constraints Tong et al. (2025).
2.2 Visual Token Pruning
@@
Visual Token Pruning aims to reduce computational overhead and improve inference efficiency by removing redundant or less informative visual tokens. A common strategy is to leverage attention scores to identify tokens for removal. For example, PruMerge Shang et al. (2024b) clusters and merges tokens in the vision encoder based on attention sparsity, while FastV Chen et al. (2024a) utilizes attention weights from the second layer of the LLM to guide pruning. SparseVLM Zhang et al. (2024b) employs cross-modal attention for text-conditioned token selection, and VisionZip Yang et al. (2025) compresses visual inputs via CLS token attention in the final vision encoder layer. FlowCut Tong et al. (2025) identifies redundancy from an information flow perspective by analyzing attention propagation. LVPruning Sun et al. (2025) uses cross-attention to assess vision token importance via interaction with language tokens, guiding token pruning. Other methods adopt calibration-based strategies, where pruning ratios and depths are determined by evaluating model behavior on a held-out set. FitPrune Ye et al. (2025) compares attention distributions before and after pruning to inform token selection, while VTW Lin et al. (2025a) shows that tokens can be safely dropped after specific layers, guided by calibration. In addition, approaches like DivPrune Alvar et al. (2025) addresses token pruning by maximizing the minimum distance between tokens. Moreover, a concurrent work Zhang et al. (2025) reformulates the token pruning problem using determinantal point processes to achieve effective pruning.
3 Token Redundancy Across Modalities
@@
Previous studies Shang et al. (2024a); Alvar et al. (2025); Li et al. (2025a) primarily decide to prune visual tokens based on the discrepancy in the number of visual and textual tokens. However, these works do not thoroughly analyze the differences in redundancy across modalities. In this work, we provide a more detailed pruning guidance by examining the redundancy of visual and textual tokens from three perspectives: attention mechanisms, semantic distribution and information repetition. Our findings reveal that visual tokens exhibit substantially higher redundancy than textual tokens. Motivated by this observation, we retain all textual tokens during the pruning process and apply pruning exclusively to visual tokens. The detailed analysis can be found in Appendix A.
4 Methodology
4.1 Problem Formulation
@@
Our pruning pipeline can be roughly divided into two stages. In the first stage, we select a subset $\mathcal{X}^{v(1)}\subset\mathcal{X}^{v}$ with size $\left|\mathcal{X}^{v(1)}\right|={N}_{1}$, where ${N}_{1}<N$, by retaining tokens with the highest semantic alignment to the text modality, thereby ensuring the preservation of tokens with strong cross-modal correlations. In the second stage, we further prune $\mathcal{X}^{v(1)}$ to obtain a more compact subset $\mathcal{X}^{v(2)}\subset\mathcal{X}^{v(1)}$ with size $\left|\mathcal{X}^{v(2)}\right|={N}_{2}$, where ${N}_{2}<{N}_{1}$, by maximizing the expected pairwise distance in the embedding space. This promotes non-redundant and information-rich representations while reducing computational overhead. We provide the experimental analysis of the stage order and the used metrics in Section 5.5 and Appendix D. Figure 1 illustrates the overview of our method.
4.2 Cross-Modal Alignment Aware Token Filtering
During training, multimodal models primarily optimize cross-entropy loss for text generation, which often leads to an implicit weakening of cross-modal alignment Covert et al. (2024). Pruning as a post-processing step for the model is crucial for cross-modal alignment. To preserve cross-modal alignment during visual token pruning, we introduce a mutual information based criterion that selects a subset of visual tokens $\mathcal{X}^{v(1)}$ which maximizes the expected mutual information with the textual tokens. Our selection objective can be formulated as follows:
$\mathcal{X}^{v(1)}=\mathop{\arg\max}_{\begin{subarray}{c}\mathcal{X}^{v^{\prime}}\subset\mathcal{X}^{v},\\ |\mathcal{X}^{v^{\prime}}|=N_{1}\end{subarray}}\mathbb{E}_{x^{t}\in\mathcal{X}^{t}}\left[I(\mathcal{X}^{v^{\prime}};x^{t})\right]$ (1)
@@ 4.2 Cross-Modal Alignment Aware Token Filtering
This score reflect the average semantic alignment strength of visual tokens within the shared embedding space. Intuitively, tokens with higher alignment scores are retained, while those with lower semantic alignment are pruned.
Although mutual information provides a theoretically grounded measure of cross-modal alignment, its precise estimation in high dimensional spaces often requires costly density modeling. We approximate mutual information using the $L_{2}$ norm, which reduces computational overhead without compromising performance. It is worth noting that our approximation is grounded in mathematical theory and does not rely on any specific model architecture. Additionally, we provide theoretical justification and experimental analysis to explain why the $L_{2}$ norm can be used for approximation (see Appendix B and D). Therefore, we can compute $\alpha_{i}$ using the following formula:
@@
$\mathcal{X}^{v(1)}=\left\{{x}_{i}^{v}\in\mathcal{X}^{v}\mid\alpha_{i}\in top_{N_{1}}(\alpha_{1},\alpha_{2},\ldots,\alpha_{N})\right\}$
This selection effectively filters out semantically misaligned visual tokens while retaining those with the strongest cross-modal correlations, thereby enhancing alignment quality and reducing computational overhead. Even for semantically sparse queries (e.g., image captioning on the COCO dataset), our method consistently achieves strong performance.
Superiority over Other Works
Some works prune visual tokens based on their attention scores with textual tokens. However, these approaches are susceptible to positional bias, often retaining spatially adjacent tokens that are highly redundant. Although both our method and these approaches leverage interactions between text and vision, our method emphasizes alignment in the semantic space rather than relying on positional correlations. Moreover, we do not rely on an additional CLIP model to remap the embeddings, as this would alter the vector space and introduce extra computational overhead. We provide a detailed comparison under different metrics in Appendix D, where the results show that the $L_{2}$ norm achieves better performance compared to other measures.
4.3 Greedy Intra-Modal Representational Maximization (RepMax)
@@
$\mathcal{X}^{v(2)}=\mathop{\arg\max}_{\begin{subarray}{c}\widetilde{\mathcal{X}}\subset\mathcal{X}^{v(1)},\\ |\widetilde{\mathcal{X}}|=N_{2}\end{subarray}}\mathbb{E}_{\begin{subarray}{c}{x}_{i}^{v},{x}_{j}^{v}\in\widetilde{\mathcal{X}},i\neq j\end{subarray}}\left[D({x}_{i}^{v},{x}_{j}^{v})\right]$ (6)
@@
$\max\limits_{\boldsymbol{\gamma}\in\{0,1\}^{N_{1}}}\sum\limits_{\begin{subarray}{c}i=1,j=1\\ i\neq j\end{subarray}}^{N_{1}}\gamma_{i}\gamma_{j}\cdot D(x_{i}^{v},x_{j}^{v}),\;\;\sum\limits_{i=1}^{N_{1}}\gamma_{i}=N_{2}$
@@
At the initial stage of the greedy algorithm, we quantify pairwise relationships among visual tokens by constructing a cosine similarity matrix $\mathcal{C}\in\mathbb{R}^{N_{1}\times N_{1}}$. Each element $\mathcal{C}_{ij}$ of the matrix $\mathcal{C}$ represents the similarity between tokens $x_{i}^{v}$ and $x_{j}^{v}$. Based on this matrix, we compute the average similarity of each token with all others. The token with the lowest average similarity is chosen as the initial seed to initialize the selected set for the first iteration ($\mathcal{S}^{(1)}$), and the remaining tokens constitute the initial remaining set ($\mathcal{R}^{(1)}$). This can be formulated as follows:
@@
$\mathcal{S}^{(1)}=\{s^{(1)}\},\;\mathcal{R}^{(1)}=[N_{1}]\setminus\{s^{(1)}\}$ (11)
@@
$\mathcal{S}^{(t)}=\mathcal{S}^{(t-1)}\cup\{s^{(t)}\},\;\mathcal{R}^{(t)}=\mathcal{R}^{(t-1)}\setminus\{s^{(t)}\}$
@@
Superiority over Other Works
We compare our method with DivPrune, which performs pruning by maximizing the minimum distance. In our experiments, we observe that outliers in the token embeddings can significantly affect DivPrune, preventing it from achieving optimal performance. In contrast, our method optimizes the expected pairwise distance, which effectively mitigates the influence of outliers, and employs a designed greedy algorithm to efficiently tackle the associated NP hard problem. The results of our comparative experiments demonstrate a substantial performance improvement over DivPrune. Furthermore, our method maintains strong performance when compared to an extended variant (DivPrune) implemented by our method (see Section 5.4).
@@
5 Experiments
5.1 Experimental Setup
Models and Baselines. We evaluate the performance of our method across several representative LMMs, including LLaVA-1.5-7B Liu et al. (2023), LLaVA-1.5-13B Liu et al. (2023), LLaVA-NEXT-7B Liu et al. (2024b), LLaVA-NEXT-Video-7B Zhang et al. (2024c) and Qwen2-VL-7B Wang et al. (2024). These models span a range of parameter scales and encompass both image and video understanding tasks. We evaluate our method alongside several widely adopted baselines, including ToMe Bolya et al. (2023), FastV Chen et al. (2024a), SparseVLM (SpVLM) Zhang et al. (2024b), PDrop Xing et al. (2024), VisionZip Yang et al. (2025) and DivPrune Alvar et al. (2025). We additionally include the VTW Lin et al. (2025b) method in our efficiency evaluation. All baselines are assessed under consistent experimental settings.
Datasets and Metrics. We evaluate our method on diverse multimodal benchmarks covering various tasks. For ChartQA Masry et al. (2022), we report the relaxed score. Image captioning quality and diversity are measured by CIDEr on COCO Sharma et al. (2018) and NoCaps Agrawal et al. (2019). MMBenchEN (MMB) Liu et al. (2024d) uses a GPT-based score. Perceptual understanding is assessed via the perception score on MME Fu et al. (2023). Accuracy evaluates general reasoning on MMU Zheng et al. (2025) and text recognition on OCRBench Liu et al. (2024e). Exact match accuracy measures question answering on VQAOK Marino et al. (2019) and VQATEXT Singh et al. (2019). POPE Li et al. (2023b) is evaluated with precision for positional understanding. This comprehensive protocol ensures thorough assessment of generalization across multimodal tasks.
Implementation Details. Our experiments are conducted using the PyTorch framework Paszke et al. (2019) and the Hugging Face Transformers library Wolf et al. (2020). We utilize an NVIDIA H100 GPU with 80GB of memory. We set the parameter $N_{1}=\left\lfloor 0.8N\right\rceil$ by default. In both the ablation study and case study, we fix the final number of retained tokens to 64. The temperature, prompt and preprocessing in our experiments all follow the default settings of LMMs-Eval.